OpenStudy (anonymous):

solve by the elimination method3x+4y=4 6x+8y=-8

OpenStudy (anonymous):

Divide the second equation by 2 it will make your calculations easier

OpenStudy (anonymous):

No solution to this system

OpenStudy (anonymous):

You will want to multiply the top one by -2 (in order to get one of the terms to cancel out. You end up with: -6x-8y=-8 6x+8y=-8 Add the two to get 0 = -16 No solution

OpenStudy (anonymous):

3x + 4y = 4 3x + 4y = -4 When you subtract equation (1) from equation (2) You have the same terms on left hand side so you get: 0 = 8...This is not true therefore your system has no solution